# Stochastic Re-weighted Gradient Descent via Distributionally Robust Optimization

### v4

@@ 3.1 Distributionally Robust Optimization
Consider a general learning problem where we are given $n$ i.i.d samples $\{z_{i}\}_{i=1}^{n}$ drawn from some unknown distribution $P_{\text{data}}$. Let $\widehat{P}_{\text{data}}$ be the empirical distribution over these samples. Our ideal goal is to find a model $\theta\in\Theta$ that minimizes the population risk: $R(\theta)\coloneqq\mathbb{E}_{P_{\text{data}}}[\ell(z;\theta)]$. Here $\ell(z;\theta)$ is the loss of $z$ under model $\theta$. Since $P_{\text{data}}$ is typically unknown, a standard practice in ML/AI is to minimize the empirical risk, which is defined as $\widehat{R}_{n}(\theta)\coloneqq\mathbb{E}_{\widehat{P}_{\text{data}}}[\ell(z;\theta)]=\frac{1}{n}\sum_{i=1}^{n}\ell(z_{i};\theta)$.
@@
$R_{D}(\theta)\coloneqq\sup_{P^{\prime}:D(P^{\prime}||P_{\text{data}})\leq\rho}\mathbb{E}_{P^{\prime}}[\ell(z,\theta)],$
$\quad\widehat{R}_{D,n}(\theta)\coloneqq\sup_{P^{\prime}:D(P^{\prime}||\widehat{P}_{\text{data}})\leq\rho}\mathbb{E}_{P^{\prime}}[\ell(z,\theta)].$
@@
Together with Equation 2, the above result shows that the empirical DRO risk $\widehat{R}_{D,n}(\theta)$ is an upper bound of the population risk $R(\theta)$ at any $\theta\in\Theta$. Consequently minimizing $\widehat{R}_{D,n}(\theta)$ achieves a better bias-variance trade-off than ERM, and leads to models with good generalization guarantees.
@@ 3.2 Stochastic Re-weighted Gradient Descent (RGD)
The above discussion motivates the use of DRO risk for learning models, especially in high variance and/or low sample regime. We now derive our RGD algorithm as a technique to minimize the empirical DRO risk $\widehat{R}_{D,n}$, and to learn models with better generalization guarantees than ERM. Specifically, we consider KL divergence-based DRO, where one adds perturbations to create data distributions that are close to the original data distribution in the KL divergence metric, and learn a model with best performance over all possible perturbations. The following proposition derives the dual representation of the KL divergence-DRO objective, the proof of which can be found in Appendix B.1.
@@
[…]
@@
7:   Update model parameters: $\theta_{t+1}\leftarrow\Pi_{\Theta}(\theta_{t}-\eta_{t}v_{t})$
@@
[…]
@@
Weight Clipping. In our experiments, when computing per-sample weights, we clip the loss $\ell$ at some constant $\tau>0$; that is, we use $g(u)=e^{\gamma\min(u,\tau)}$. We observed this clipping to help stabilize the training in the presence of outliers. We note that this is different from other common techniques, such as loss, gradient clipping, which are used to make the learning process robust to outliers (Yang et al., 2010; Catoni & Giulini, 2018; Menon et al., 2019; Koloskova et al., 2023). We plan to investigate the robustness properties of weight clipping in more detail in future work. In our experiments, we choose the scale parameter $\gamma=1/(\tau+1)$. Even with this fixed choice of $\gamma$, our algorithm provides a significant boost in performance over vanilla optimization techniques (see Section 4.1 for ablations on our design choices).
Choice of divergences. One could rely on other divergences instead of the KL-divergence we used in the above derivation. From a theoretical perspective, DRO with many $f$-divergences (KL, reverse KL, chi-squared etc.) provides an upper bound on the true population risk (Lam, 2016; Duchi et al., 2021). For example, using $\chi^{2}$-divergence ($f(x)=(x-1)^{2}$) gives us the following re-weighting function for positive loss functions: $g(u)=u+\tau$. Using reverse KL-divergence ($f(x)=-\log{x}$) gives us the following re-weighting function $g(u)=\frac{1}{\tau-u}$ for some appropriate choice of $\tau$ (see Appendix B.3 for more details). Observe that the reverse KL-divergence based weighting function is much more aggressive in up-weighting high loss points than KL-divergence based weighting function. In the sequel, we denote the approach with $g(u)=u+\tau$ as RGD-$\chi^{2}$, $g(u)=\frac{1}{\tau-u}$ as RGD-RevKL, and $g(u)=e^{\min(u,\tau)/(\tau+1)}$ as RGD. While all the three techniques provided a performance boost over ERM (see Appendix D), KL-divergence based reweighting has better performance than reverse KL, chi-squared divergences. However, the current theoretical understanding of DRO doesn’t explain this nuanced behavior and we believe that this is an interesting direction for future research. From a practitioner perspective, the choice of $f$-divergence could be treated as a hyper-parameter which could be tuned using cross-validation.
@@ 4 Experiments
In this section, we first present ablations on various design choices in our algorithm. Next, we present empirical evidence showing that RGD outperforms ABSGD (Qi et al., 2020), TERM (Li et al., 2020), two state-of-the-art algorithms for optimizing KL-DRO. Finally, we present large scale experiments showing that RGD can be widely applied across tasks such as supervised learning, meta learning to boost the generalization performance of existing learning algorithms. Due to space constraints, details regarding hyperparameter tuning are relegated to Appendix Reproducibility Statement. Additional experiments on class imbalanced classification, and large-scale tasks such as EfficientNet finetuning, LLM pretraining (miniGPT) are presented in Appendix F.
@@ 4.1 Ablation Studies
Choice of scale parameter ($\gamma$). Recall, RGD uses $\gamma=\frac{1}{\tau+1}$, where $\tau$ is the clipping level. To understand the utility of this choice, we vary $\gamma$. To be precise, we set $\tau=1$ and get accuracy numbers of RGD with various values of $\gamma$. Figure 1 (light blue line) presents the results from this experiment. For each value of $\gamma$, we report the accuracy obtained using the best learning rate identified using hold-out set validation. It can be seen that RGD is fairly robust to the choice of $\gamma$.
@@
Importance of clipping. We now study the importance of clipping. To this end, we replace the proposed reweighting function in RGD with $g(u)=e^{\gamma u}$. Figure 1 (dark blue line) presents the results from this experiment, for various values of $\gamma$. It can be seen that the best accuracy without clipping is $1\%$ off compared to RGD with clipping (with $\tau=1$). We believe this performance drop primarily happens because of the high weights given to the outliers. This demonstrates the importance of clipping. Next, we vary the clipping factor $\tau$. It is evident that as $\tau$ increases, the performance of RGD drops and approaches the performance of no clipping. This shows the importance of setting an appropriate $\tau$.
@@
Default 81.33 89.62 87.93 90.63 89.55 67.19 54.53 80.11 78.1 71.31 98.01
RGD (Ours) 83.06 91.06 90.35 91.78 88.28 71.48 58.56 82.05 79.11 $\pm$ 0.12 71.51 98.91
4.2 Comparison with existing KL-DRO optimization techniques
In this section, we present experimental results on ImageNet classification with ViT-S model to showcase the efficacy of RGD over ABSGD, and TERM (see Table 3). We were unable to compare with the recent SCDRO algorithm of Qi et al. (2022) due to technical difficulties in implementing it in JAX (Bradbury et al., 2018) (the main difficulty arises from syncing values of parameters across various devices). Furthermore, the algorithm is significantly more complex with many hyperparameters. Table 3 shows that ABSGD, TERM with extensive tuning of three hyperparameters (exponential scale, learning rate , and moving average parameter) achieve similar performance as baseline ERM. In contrast, RGD outperforms the baseline ERM by $1.1\%$ with minimal tuning of hyperparameters, highlighting its efficacy. One of the reasons for this performance difference between RGD, and ABSGD, TERM is the weight clipping we perform in our algorithm, which guards it from outliers. We note that SCDRO doesn’t perform weight clipping and could potentially suffer from a similar performance drop as ABSGD, TERM. Additional details about the experiment can be found in Appendix F.3.
Next, we compare RGD with ABSGD, TERM for the problem of class imbalanced classification. For this experiment, we consider the long-tailed CIFAR-10, CIFAR-100 datasets (Cui et al., 2019). Table 4 presents the results from this experiment. It can be seen that RGD outperforms both ABSGD and TERM by $>1\%$ on average. Additional details about this experiment, including comparison with specialized techniques for class imbalanced classification such as class-balanced loss (Cui et al., 2019) and focal loss (Lin et al., 2017), can be found in Appendix F.2.
Algorithm ERM TERM ABSGD RGD
Accuracy 78.1 78.4 78.28 79.11 $\pm$ 0.12
Table 4: Comparison of RGD with existing KL-DRO optimization techniques for class imbalanced classification. The numbers represent test accuracy on Long-Tailed CIFAR-10, and CIFAR-100 datasets using ResNet-32.
Dataset CIFAR-10 CIFAR-100
Loss / Imbalance Factor 100 10 Avg. 100 10 Avg.
Cross Entropy (CE)
Default 70.36 86.39 78.38 38.32 55.71 47.02
TERM Li et al. (2020) 72.19 87.56 79.88 39.7 57.37 48.57
ABSGD Qi et al. (2020) 72.43 87.93 80.18 39.77 57.44 48.61
RGD (Ours) 73.75 88.00 80.88 41.89 58.5 50.20
4.3 Supervised Learning
This section studies our approach when applied on standard supervised learning tasks such as BERT finetuning on GLUE benchmark, and Imagenet-1K classification. We use a base model of ViT-S for the latter task. Table 2 depicts our results from this experiment. On GLUE tasks, our RGD algorithm outperforms the baseline by +1.94%. On Imagenet-1K, we show a +1.01% improvement over baseline with the off-the-hat addition of the RGD reweighing and no additional complexity in terms of compute: memory and time. Furthermore, we also experiment with pre-training of the BERT-base model. We use the BooksCorpus (800M words) (Zhu et al., 2015) and English Wikipedia (2,500M words) as our pre-training corpus. We trained the Bert-base model for 450K steps, and tuned the learning rate (lr) for baseline and lr, clipping factor for RGD. We report both the MLM (Masked Language Model) accuracy and NSP (Next Sequence Prediction) accuracy comparisons of RGD vs Default (ERM). It can be seen that our approach boosts the MLM accuracy and NSP accuracy by +0.2% and +0.9% respectively (see Table 2). Additional experiments on EfficientNet and miniGPT are discussed in Appendix F.4.

### v5

@@ 3.1 Distributionally Robust Optimization
Consider a general learning problem where we are given $n$ i.i.d samples $\{z_{i}\}_{i=1}^{n}$ drawn from some unknown distribution $P_{\text{data}}$. Let $\widehat{P}_{\text{data}}$ be the empirical distribution over these samples. Our ideal goal is to find a model $\theta\in\Theta$ that minimizes the population risk: $R(\theta)\coloneqq\mathbb{E}_{P_{\text{data}}}[\ell(z;\theta)]$. Here $\ell(z;\theta)$ is the loss of $z$ under model $\theta$. Since $P_{\text{data}}$ is typically unknown, a standard practice in ML/AI is to minimize the empirical risk, which is defined as
$\widehat{R}_{n}(\theta)\coloneqq\mathbb{E}_{\widehat{P}_{\text{data}}}[\ell(z;\theta)]=\frac{1}{n}\sum_{i=1}^{n}\ell(z_{i};\theta).$
@@
$R_{D}(\theta)\coloneqq\sup_{P^{\prime}:D(P^{\prime}||P_{\text{data}})\leq\rho}\mathbb{E}_{P^{\prime}}[\ell(z,\theta)],\quad\widehat{R}_{D,n}(\theta)\coloneqq\sup_{P^{\prime}:D(P^{\prime}||\widehat{P}_{\text{data}})\leq\rho}\mathbb{E}_{P^{\prime}}[\ell(z,\theta)].$
@@
Ignoring the lower order terms (i.e., $1/n$ terms), the above equation, together with Equation 2, shows that the empirical DRO risk $\widehat{R}_{D,n}(\theta)$ is an upper bound of the population risk $R(\theta)$ at any $\theta\in\Theta$. Furthermore, it can be seen that empirical DRO is equal to the empirical risk plus a variance term (modulo the lower order term). This variance term acts as a regularizer during the optimization of empirical DRO and leads to models with smaller variance, and with good generalization guarantees.
@@ 3.2 Stochastic Re-weighted Gradient Descent (RGD)
The above discussion motivates the use of DRO risk for learning models, especially in high variance and/or low sample regime. We now derive our RGD algorithm as a technique to minimize the empirical DRO risk $\widehat{R}_{D,n}$, and to learn models with better generalization guarantees than ERM. Specifically, we consider KL divergence-based DRO, where one adds perturbations to create data distributions that are close to the original data distribution in the KL divergence metric, and learn a model with best performance over all possible perturbations. The following proposition derives the equivalent dual representation of the KL divergence-DRO objective, the proof of which can be found in Appendix B.1.
@@
(Shapiro, 2017) Consider DRO with KL-divergence-based uncertainty set. Then $\min_{\theta\in\Theta}\widehat{R}_{D,n}$ can be rewritten as:
$\min_{\theta\in\Theta}\frac{1}{\gamma}\log{\mathbb{E}_{\widehat{P}_{\text{data}}}[e^{\gamma\ell(z;\theta)}]},$
for some constant $\gamma>0$ that is independent of $\theta$.
@@
7:   Update model parameters:
$\theta_{t+1}\leftarrow\Pi_{\Theta}(\theta_{t}-\eta_{t}v_{t})$
@@
Assume that $\Theta$ is a convex and compact set. For all data points $z$ suppose that $\ell(z;\cdot)$ is convex, continuously differentiable and bounded in the range $[-M,M]$, and $\nabla\ell(z;\cdot)$ is uniformly bounded over the set $\Theta$. Let the step-size sequence $(\eta_{t})_{t}$ be such that $\eta_{t}=\frac{C}{\sqrt{t}}$ $\forall$ $1\leq t\leq T$ or $\eta_{t}=\frac{C}{\sqrt{T}}$ $\forall$ $1\leq t\leq T$. Then the sub-optimality gap satisfies:
$\mathbb{E}_{\theta_{T}}\frac{1}{\gamma}\log{\mathbb{E}_{\widehat{P}_{\text{data}}}[e^{\gamma\ell(z;\theta_{T})}]}-\min_{\theta\in\Theta}\frac{1}{\gamma}\log{\mathbb{E}_{\widehat{P}_{\text{data}}}[e^{\gamma\ell(z;\theta)}]}=O\left(\frac{\log T}{\sqrt{T}}\right).$
@@
Weight Clipping.
In our experiments, when computing per-sample weights, we clip the loss $\ell$ at some constant $\tau>0$; that is, we use $g(u)=e^{\gamma\min(u,\tau)}$. We observed this clipping to help stabilize the training in the presence of outliers (see Section 4.1 for empirical evidence). We note that this is different from other common techniques, such as loss, gradient clipping, which are used to make the learning process robust to outliers in empirical risk minimization (Yang et al., 2010; Catoni & Giulini, 2018; Menon et al., 2019; Koloskova et al., 2023). We plan to investigate the robustness properties of weight clipping in more detail in future work. In our experiments, we choose the scale parameter $\gamma=1/(\tau+1)$. Even with this fixed choice of $\gamma$, our algorithm provides a significant boost in performance over vanilla optimization techniques (see Section 4.1 for ablations on our design choices).
Choice of divergences.
One could rely on other divergences instead of the KL-divergence we used in the above derivation. From a theoretical perspective, DRO with many $f$-divergences (KL, reverse KL, chi-squared etc.) provides an upper bound on the true population risk (Lam, 2016; Duchi et al., 2021b). For example, using $\chi^{2}$-divergence ($f(x)=(x-1)^{2}$) gives us the following re-weighting function for positive loss functions: $g(u)=u+\tau$. Using reverse KL-divergence ($f(x)=-\log{x}$) gives us the following re-weighting function $g(u)=\frac{1}{\tau-u}$ for some appropriate choice of $\tau$ (see Appendix B.3 for more details). Observe that the reverse KL-divergence based weighting function is much more aggressive in up-weighting high loss points than KL-divergence based weighting function. In the sequel, we denote the approach with $g(u)=u+\tau$ as RGD-$\chi^{2}$, $g(u)=\frac{1}{\tau-u}$ as RGD-RevKL, and $g(u)=e^{\min(u,\tau)/(\tau+1)}$ as RGD. While all the three techniques provided a performance boost over ERM (see Appendix C), KL-divergence based reweighting has better performance than reverse KL, chi-squared divergences. However, the current theoretical understanding of DRO doesn’t explain this nuanced behavior and we believe that this is an interesting direction for future research. From a practitioner perspective, the choice of $f$-divergence could be treated as a hyper-parameter which could be tuned using cross-validation.
@@ 4 Experiments
In this section, we first present ablations on various design choices in our algorithm. Next, we present empirical evidence showing that RGD outperforms ABSGD (Qi et al., 2023b), TERM (Li et al., 2021), two state-of-the-art algorithms for optimizing KL-DRO. Finally, we present large scale experiments showing that RGD can be widely applied across tasks such as supervised learning, meta learning to boost the generalization performance of existing learning algorithms. Details regarding hyperparameter tuning are relegated to Appendix Reproducibility Statement. Additional experiments on class imbalanced classification, and large-scale tasks such as miniGPT pre-training, EfficientNet finetuning are presented in Appendix D.
To ensure fair comparisons, we integrated RGD into existing baseline codebase for each of the experiments, maintaining the same optimizer (primarily Adam for most experiments, SGD for CIFAR experiments) for both baseline and RGD versions. RGD introduces only one additional hyperparameter, the clipping factor ($\tau$). The optimizer, weight decay, batch sizes, etc. were kept constant with the baseline across all our experiments.
@@ 4.1 Ablation Studies
Recall, RGD uses $\gamma=\frac{1}{\tau+1}$, where $\tau$ is the clipping level. To understand the utility of this choice, we vary $\gamma$. To be precise, we set $\tau=1$ and get accuracy numbers of RGD with various values of $\gamma$. Figure 1 (light blue line with stars) presents the results from this experiment. For each value of $\gamma$, we report the accuracy obtained using the best learning rate identified using hold-out set validation. It can be seen that RGD is fairly robust to the choice of $\gamma$.
Importance of clipping.
@@
We now study the importance of clipping. To this end, we replace the proposed reweighting function in RGD with $g(u)=e^{\gamma u}$. Figure 1 (dark blue line with squares) presents the results from this experiment, for various values of $\gamma$. It can be seen that the best accuracy without clipping is $1\%$ off compared to RGD with clipping (with $\tau=1$). We believe this performance drop primarily happens because of the high weights given to the outliers. This demonstrates the importance of clipping. Next, we vary the clipping factor $\tau$. It is evident that as $\tau$ increases, the performance of RGD drops and approaches the performance of no clipping. This shows the importance of setting an appropriate $\tau$.
To further demonstrate that our weight clipping can effectively handle benign outliers, we perform the following experiment. We randomly flip the labels in CIFAR-10, CIFAR-100 datasets (we vary the proportion of flips from $0\%$ to $40\%$) and compare the performance of RGD with state-of-the-art KL-DRO optimization technique TERM (Li et al., 2021). The results from this experiment are reported in Table 2. Note that TERM, which does not account for outliers, performs poorly in this experiment, as it upweights the corrupted/mislabeled points. Whereas, RGD which clips the weights of outliers, achieves the best performance.
Table 2: Results on CIFAR-10 and CIFAR-100 dataset with corrupted labels.
Dataset CIFAR-10 CIFAR-100
Loss 0% 20% 40% Avg. 0% 20% 40% Avg.
Default (Cross Entropy) 92.89 $\pm$ 0.32 76.83 $\pm$ 2.30 70.77 $\pm$ 2.31 80.16 70.50 $\pm$ 0.12 50.86 $\pm$ 0.27 43.01 $\pm$ 1.16 54.79
TERM (Li et al., 2021) 92.90 $\pm$ 0.09 58.7 $\pm$ 39.76 73.17 $\pm$ 31.59 74.92 70.42 $\pm$ 0.41 63.85 $\pm$ 1.40 46.59 $\pm$ 22.80 60.29
RGD (Ours) 93.04 $\pm$ 0.24 90.69 $\pm$ 0.24 88.90 $\pm$ 0.15 90.88 71.06 $\pm$ 0.22 64.61 $\pm$ 0.39 57.17 $\pm$ 0.80 64.28
4.2 Comparison with existing KL-DRO optimization techniques
In this section, we present experimental results on ImageNet classification with ViT-S model to showcase the efficacy of RGD over state-of-the-art KL-DRO optimization techniques ABSGD, and TERM (see Table 3). We were unable to compare with the recent SCDRO algorithm of Qi et al. (2023a) due to technical difficulties with implementing it in JAX (Bradbury et al., 2018) (the main difficulty arises from syncing values of parameters across various devices). Furthermore, the algorithm is significantly more complex with many hyperparameters. Table 3 shows that ABSGD, TERM with extensive tuning of three hyperparameters (exponential scale, learning rate , and moving average parameter) achieve similar performance as baseline ERM. In contrast, RGD outperforms the baseline ERM by $1.1\%$ with minimal tuning of hyperparameters, highlighting its efficacy. One of the reasons for this performance difference between RGD, and ABSGD, TERM is the weight clipping we perform in our algorithm, which guards it from outliers. We note that SCDRO doesn’t perform weight clipping and could potentially suffer from a similar performance drop as ABSGD, TERM. Additional details about the experiment can be found in Appendix D.1.1.
Next, we compare RGD with ABSGD, TERM for the problem of class imbalanced classification. For this experiment, we consider the long-tailed CIFAR-10, CIFAR-100 datasets (Cui et al., 2019). Table 4 presents the results from this experiment. It can be seen that RGD outperforms both ABSGD and TERM by $>1\%$ on average. Additional details about this experiment, including comparison with specialized techniques for class imbalanced classification such as class-balanced loss (Cui et al., 2019) and focal loss (Lin et al., 2017), can be found in Appendix D.2.
Algorithm ERM TERM ABSGD RGD
Accuracy 78.44 $\pm$ 0.35 78.03 $\pm$ 0.27 77.12 $\pm$ 0.82 79.11 $\pm$ 0.12
Table 4: Comparison of RGD with existing KL-DRO optimization techniques for class imbalanced classification. The numbers represent test accuracy on Long-Tailed CIFAR-10, and CIFAR-100 datasets using ResNet-32. Moreover, we perform hypothesis tests to confirm that these results are statistically significant with a p-value that is less than 0.05.
Dataset CIFAR-10 CIFAR-100
Loss / Imbalance Factor 100 10 Avg. 100 10 Avg.
Cross Entropy (CE)
Default 71.75 $\pm$ 0.75 87.64 $\pm$ 0.45 79.7 38.35 $\pm$ 38.35 56.91 $\pm$ 0.41 47.63
TERM Li et al. (2021) 72.20 $\pm$ 0.18 87.52 $\pm$ 0.07 79.86 39.75 $\pm$ 0.04 57.91 $\pm$ 0.39 48.83
ABSGD Qi et al. (2023b) 72.43 $\pm$ 0.31 87.93 $\pm$ 0.25 80.18 39.77 $\pm$ 0.34 57.44$\pm$ 0.25 48.61
RGD (Ours) 73.71 $\pm$ 0.04 87.78 $\pm$ 0.15 80.75 41.47 $\pm$ 0.30 58.49 $\pm$ 0.04 49.98
4.3 Supervised Learning
This section studies our approach when applied on standard supervised learning tasks such as BERT finetuning on GLUE benchmark, and Imagenet-1K classification. We use a base model of ViT-S for the latter task. Table 5 depicts our results from this experiment. On GLUE tasks, our RGD algorithm outperforms the baseline by +1.94% with a standard deviation of 0.42%. Furthermore, we perform hypothesis testing to condit that these results are statistically significant with a p-value that is less than 0.05. On Imagenet-1K, we show a +1.01% improvement over baseline with the off-the-hat addition of the RGD reweighing and no additional complexity in terms of compute: memory and time.
Furthermore, we also experiment with pre-training of the BERT-base model. We use the BooksCorpus (800M words) (Zhu et al., 2015) and English Wikipedia (2,500M words) as our pre-training corpus. We trained the Bert-base model for 450K steps, and tuned the learning rate (lr) for baseline, and lr, clipping factor for RGD. We report both the MLM (Masked Language Model) accuracy and NSP (Next Sequence Prediction) accuracy comparisons of RGD vs Default (ERM). It can be seen that our approach boosts the MLM accuracy and NSP accuracy by +0.2% and +0.9% respectively (see Table 5). Furthermore, through hypothesis testing, we show that the results are statistically significant with a p-value of 0.05. Additional experiments on EfficientNet fine-tuning, DeiT model (Touvron et al., 2021) for ImageNet-1K classification, and miniGPT (Zhu et al., 2024) pre-training are discussed in Appendix D.5.
@@ 4.3 Supervised Learning
Default 81.33 89.62 87.93 90.63 89.55 67.19 54.53 80.11 78.44 $\pm$ 0.35 71.31 98.01
RGD (Ours) 83.06 91.06 90.35 91.78 88.28 71.48 58.56 82.05 79.11 $\pm$ 0.12 71.35 $\pm$ 0.11 98.92 $\pm$ 0.09
4.4 Tabular Classification